### v1

# FlowHFT: Flow Policy Induced Optimal High-Frequency Trading under Diverse Market Conditions

Abstract
High-frequency trading (HFT) is an investing strategy that continuously monitors market states and places bid and ask orders at millisecond speeds. Traditional HFT approaches fit models with historical data and assume that future market states follow similar patterns. This limits the effectiveness of any single model to the specific conditions it was trained for. Additionally, these models achieve optimal solutions only under specific market conditions, such as assumptions about stock price’s stochastic process, stable order flow, and the absence of sudden volatility. Real-world markets, however, are dynamic, diverse, and frequently volatile. To address these challenges, we propose the FlowHFT, a novel imitation learning framework based on flow matching policy. FlowHFT simultaneously learns strategies from numerous expert models, each proficient in particular market scenarios. As a result, our framework can adaptively adjust investment decisions according to the prevailing market state. Furthermore, FlowHFT incorporates a grid-search fine-tuning mechanism. This allows it to refine strategies and achieve superior performance even in complex or extreme market scenarios where expert strategies may be suboptimal. We test FlowHFT in multiple market environments. We first show that flow matching policy is applicable in stochastic market environments, thus enabling FlowHFT to learn trading strategies under different market conditions. Notably, our single framework consistently achieves performance superior to the best expert for each market condition.
1 Introduction
Market making (MM) is a trading strategy where participants place both buy and sell orders to profit from difference in the bid-ask spread. By providing liquidity through these actions, market makers also enhance market efficiency  madhavan2000market ; harris2002trading . High-frequency trading (HFT) is one type of market making that executes orders at high speeds, often measured in milliseconds or microseconds. This rapid execution allows traders to earn profit on very small price changes within short timeframes.  avellaneda2008high ; cartea2014buy ; ait2017high .
Traditional HFT strategies, such as the Avellaneda-Stoikov (AS) model  avellaneda2008high and the Guéant-Lehalle-Fernandez-Tapia (GLFT) model  gueant2013dealing , rely on historical market data to calibrate their parameters. These calibrated models are used to place orders in subsequent market. However, the effectiveness of traditional HFT models is constrained by two key limitations. If future market conditions diverge substantially from the historical period used for calibration, the optimized parameters may become ineffective and lead to suboptimal performance. Moreover, these models are often built on optimization frameworks requiring strict assumptions about market behavior (e.g., specific stock price processes, stable order flow, no sudden jumps) that reflect idealized rather than real-world scenarios. Consequently, when actual market conditions violate these assumptions—a common occurrence—these traditional models struggle to adapt and tend to produce suboptimal trading outcomes.
Reinforcement learning (RL) offers another approach to high-frequency trading, framing the problem as an agent (the trading algorithm) interacting with an environment (the market). The agent observes the market state, takes actions and receives rewards based on metrics like profit and loss. Through this interaction, the RL aims to learn a policy—a strategy—that maximizes cumulative rewards. However, RL often focuses on optimizing only single-step actions, which can lead to compounding error  zheng2024reinforcement ; ganesh2018deep  kober2013reinforcement . for instance, slightly suboptimal actions at each step might accumulate, resulting in a large unwanted position. Effective HFT necessitates a long-term view, because profitability typically relies on accumulating numerous small gains from individual steps. Moreover, each action taken interacts with the market environment, thereby impacting the feasibility and profitability of subsequent moves.
To address the challenges associated with traditional HFT methods, we propose FlowHFT, a novel imitation learning framework based on flow matching policy  sohl2015deep ; ho2020denoising ; song2020score  janner2022planning ; wang2022diffusion . This framework consists of two primary components. The first component utilizes imitation learning to develop a pre-trained model for observing the market state and generate corresponding trading actions. We simulate diverse market scenarios and evaluate various traditional models within each. For every scenario, the model achieving the best performance is designated as the ’expert’. FlowHFT then learns by imitating the actions of these designated experts across all scenarios, effectively integrating the strongest known strategy for each market condition into one single, comprehensive model. Imitation learning is useful because these models can indeed generate some optimal solutions in some scenarios. Even in those scenarios where they only generate suboptimal solutions, FlowHFT can still learn their strategies and improve on them subsequently. The second component uses a grid-search mechanism to fine-tune the action proposed by the pre-trained model. The initial action proposed by the pre-trained model, refined by the adjustment from the fine-tuning stage, constitutes the final action.
To the best of our knowledge, this study represents the first application of flow matching policy to a financial stochastic control problem. The results show that the FlowHFT framework integrates knowledge from multiple expert strategies into a single, adaptive model. When market scenarios shift, the framework demonstrates adaptability by generating actions appropriate to the prevailing conditions, effectively leveraging patterns learned from relevant expert demonstrations. Additionally, FlowHFT is designed for rapid inference, enabling the generation of trading actions at millisecond speeds for HFT tasks. Furthermore, the grid search mechanism refine these generated actions. Experimental results indicate that FlowHFT consistently achieves superior performance compared to the best results obtained from baseline models individually optimized for each specific market scenario.
The results indicate that our model exhibits significant robustness to market conditions involving sudden price jumps. We attribute this resilience primarily to our method’s output: it generates sequences of actions over a planning horizon, rather than isolated single-step decisions. By producing action sequences, the model inherently considers near-term trajectory and allows for smoother adjustments, effectively taking into account the broader consequences of initial moves. This sequential planning perspective helps mitigate the compounding error, aligning actions more effectively over time and thus enhancing strategic stability and performance during volatile events like jumps.
2 Preliminaries
In this section, we formulate the high frequency trading task mathematically. High-frequency trading utilizes automated algorithms for order execution at extremely high speeds, often operating within millisecond or microsecond timescales. The objective is to continuously place buy (bid) and sell (ask) limit orders for a specific financial instrument, while aiming to profit from the bid-ask spread.
We can model the HFT market making task as a stochastic control process avellaneda2008high ; gueant2013dealing defined over a set of discrete time steps $\mathbb{T}=\{0,1,\dots,T\}$. The observation state space $\mathcal{O}$ captures relevant information at time $t\in\mathbb{T}$. An observation state $O_{t}\in\mathcal{O}$ typically includes observable market information $L_{t}$ (often derived from the Limit Order Book, LOB, including stock prices and bid-ask spread) and the agent’s information $Z_{t}$, including balance, current inventory level , along with the time $t$ itself, such that $\mathcal{O}\subseteq\mathcal{L}\times\mathbb{Z}\times\mathbb{T}$, where $\mathcal{L}$, $\mathbb{Z}$ is the space of market information and agent information, respectively. We generally assume the state $S_{t}$ satisfies the Markov property.
Within this framework, the market maker chooses an action $A_{t}$ from the action space $\mathcal{A}$. A common action involves setting the agent’s bid and ask quotes, often parameterized by spreads $(\delta^{b}_{t},\delta^{a}_{t})$ relative to a reference price $p^{ref}_{t}$ derived from the market state $L_{t}$. The system’s evolution is governed by stochastic transition probabilities $P(O_{t+1}|O_{t},A_{t})$, defining the likelihood of moving to the next observation state $O_{t+1}=(L_{t+1},I_{t+1},t+1)$ given the current state and action. This transition reflects changes in both market conditions $L_{t+1}$ and the agent’s inventory $I_{t+1}$ due to order executions and market activity. The agent seeks an optimal policy $\pi:\mathcal{O}\to\mathcal{A}$, a rule mapping states to actions ($A_{t}=\pi(O_{t})$), that maximizes a defined objective function, $J(\pi)$. A standard objective is to maximize the expected final value, often combining terminal cash wealth $W_{T}$ with a penalty $\phi(I_{T})$ for inventory risk at the horizon $T$. The optimization problem is thus:
$\max_{\pi}J(\pi)=\max_{\pi}\mathbb{E}^{\pi}\left[W_{T}-\phi(I_{T})\mid O_{0}\right]$ (1)
Where, $\mathbb{E}^{\pi}[\cdot]$ denotes the expectation under policy $\pi$ starting from the initial observation state $O_{0}=(L_{0},I_{0},0)$. Solving for the optimal policy $\pi^{*}=\arg\max_{\pi}J(\pi)$ requires dynamically balancing profitability against inventory and adverse selection risks in a complex, stochastic environment.
3 Architecture of FlowHFT
Figure 1: FlowHFT framework
3.1 Backbone of FlowHFT: Flow Matching Policy
The core of FlowHFT is a conditional flow matching policy lipman2022flow , denoted $\pi_{\theta}(A_{t+1}\mid O_{t})$, which learns to map sequences of market observations $O_{t}$ to trading actions $A_{t+1}$. This policy is trained via imitation learning from expert demonstrations. We draw inspiration from recent successes of flow matching in domains like robotics, where these models effectively learn complex, continuous control policies from expert data by mapping high-dimensional inputs to sophisticated, sequential actions.
Our objective is to train a neural network $v_{\theta}(a,t\mid O_{t})$ to parameterize a conditional vector field. This vector field defines a probability flow that transports samples from a simple prior distribution $p_{\mathrm{prior}}(a_{0})$ (e.g., Gaussian noise) to the distribution of expert trading action sequences $p_{\mathrm{expert}}(a_{E}\mid O_{t})$, conditioned on the observed market state $O_{t}$. Following seminal work on flow matching, we define a time-dependent probability path $p_{t}(a\mid a_{0},a_{E},O_{t})$ that interpolates between an initial random action sequence $a_{0}$ and an expert action sequence $a_{E}$. A simple linear interpolation is given by:
$a_{t}=(1-t)\,a_{0}+t\,a_{E},\quad u_{t}(a_{t}\mid a_{0},a_{E})=a_{E}-a_{0}.$ (2)
The training procedure (Algorithm 1) minimizes the discrepancy between the network’s predicted vector field $v_{\theta}(a_{t},t\mid O_{t})$ and the target vector field $u_{t}(a_{t}\mid a_{0},a_{E})$. Specifically, given a market state $O_{t}$, the model samples a random initial action sequence $a_{0}\sim p_{\mathrm{prior}}$ and a corresponding expert action sequence $a_{E}\sim p_{\mathrm{expert}}(\cdot\mid O_{t})$. It then samples a time $t\sim\mathcal{U}(0,1)$ and forms an interpolated action sequence $a_{t}$ using the defined path. The network is trained by minimizing the following Flow Matching (FM) loss:
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t,O_{t},a_{0},a_{E},a_{t}\sim p_{t}(a\mid a_{0},a_{E},O_{t})}\left[\|v_{\theta}(a_{t},t\mid O_{t})-u_{t}(a_{t}\mid a_{0},a_{E})\|^{2}\right].$ (3)
This objective effectively teaches the network $v_{\theta}$ to generate the flow that transforms noise into expert-like HFT action sequences conditioned on the market state.
At inference (Algorithm 2), to generate an action sequence conditioned on a live market observation $O_{t}$, the policy starts with an action sequence $a_{0}\sim p_{\mathrm{prior}}$. It then iteratively refines this sequence by solving the ordinary differential equation: $\frac{da}{dt}=v_{\theta}(a,t\mid O_{t}),$ from $t=0$ to $t=1$. This is typically achieved using a numerical ODE solver (e.g., Euler or Heun method) over a small number of discretization steps $N$:
$a_{k+1}=a_{k}+\frac{1}{N}\;v_{\theta}(a_{k},t_{k}\mid O_{t}),\quad t_{k}=k/N,\;k=0,\dots,N-1.$ (4)
This process transforms the initial noise sequence into a coherent, market-aware sequence of high-frequency trading actions.
Parameters: Neural network $v_{\theta}(a,t|O_{t})$ with parameters $\theta$, learning rate $\eta$
Input: Expert demonstration data $D_{\text{expert}}$ containing pairs of $(O_{t},a_{E}(O_{t}))$
1
2do
3       Sample a market observation $O_{t}$ and corresponding expert action sequence $a_{E}\sim p_{\text{expert}}(\cdot|O_{t})$ from $D_{\text{expert}}$ (or generate via simulation/expert);
4       Sample an initial action sequence $a_{0}\sim p_{\text{prior}}(\cdot)$ (e.g., $a_{0}\sim\mathcal{N}(0,I)$);
5       Sample a time $t_{\text{sample}}\sim\mathcal{U}(0,1)$;
6
7      $a_{t_{\text{sample}}}\leftarrow(1-t_{\text{sample}})a_{0}+t_{\text{sample}}a_{E}$;
8       $u_{t_{\text{sample}}}\leftarrow a_{E}-a_{0}$;
9
10      $\hat{v}_{t_{\text{sample}}}\leftarrow v_{\theta}(a_{t_{\text{sample}}},t_{\text{sample}}|O_{t})$;
11       $\mathcal{L}\leftarrow\|\hat{v}_{t_{\text{sample}}}-u_{t_{\text{sample}}}\|^{2}$
12      $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
13
14until not converged;
Algorithm 1 Flow Matching Policy Training
Parameters: Trained policy network $v_{\theta}(a,t|O_{t})$, Number of discretization steps $N$
Input: Current market observation $O_{t}^{\text{current}}$
1 $a\sim p_{\text{prior}}(\cdot)$ (e.g., $a\sim\mathcal{N}(0,I)$)
2 $t_{\text{current}}\leftarrow 0$;
3 $\Delta t_{\text{step}}\leftarrow 1/N$;
4
5for $k\leftarrow 0$ to $N-1$ do
6       $v_{k}\leftarrow v_{\theta}(a,t_{\text{current}}|O_{t}^{\text{current}})$
7       $a\leftarrow a+v_{k}\cdot\Delta t_{\text{step}}$
8       $t_{\text{current}}\leftarrow t_{\text{current}}+\Delta t_{\text{step}}$;
9
10
Output: $a$
Algorithm 2 Flow Matching Policy Inference
3.2 Accelerated Inference via Shortcut Policy
While the flow matching policy described in Section 3.1 can accurately model the desired conditional distribution $p(A_{t+1}|O_{t})$, the iterative nature of ODE solving (as in Eq. (4)) with many steps $N$ might be too slow for the stringent low-latency requirements of high-frequency trading (HFT). To address this, we choose shortcut policy frans2024one , denoted $s_{\phi}(a,t,\Delta t|O_{t})$, which is specifically trained to generate high-quality action sequences in significantly fewer discretization steps. This policy is conditioned not only on the state $a$, time $t$, and observation $O_{t}$, but also on the intended integration step size $\Delta t$.
The training of this shortcut policy $s_{\phi}$ (Algorithm 3, where $s_{\theta}$ in the algorithm represents $s_{\phi}$ with parameters $\phi$) is designed to enable it to take larger, more informed discrete steps. The procedure combines two key strategies, leveraging insights from rectified flow and consistency models.
Firstly, for a portion of the training data (Algorithm 3, lines 4-7), the policy is guided by the direct target vector $x_{1}-x_{0}$ (analogous to $a_{E}-a_{0}$ in our action space context). Here, $x_{0}\sim\mathcal{N}(0,I)$ is a noise sample, $x_{1}\sim D$ is an expert sample, and $x_{t}$ is an interpolation. By setting $d=0$ (line 7) for these elements, the model $s_{\phi}(x_{t},t,0|O_{t})$ is implicitly trained to align with this direct path velocity, encouraging straighter, more efficient generation paths.
Secondly, for other batch elements (Algorithm 3, lines 8-12), the policy is trained using a self-consistency objective inspired by methods like progressive distillation. The model $s_{\phi}(x_{t},t,2\Delta t_{\text{small}}|O_{t})$ is trained to predict a velocity whose application over a larger step $2\Delta t_{\text{small}}$ matches the outcome of two consecutive smaller steps, each of size $\Delta t_{\text{small}}$. The target for this larger step is derived from stop-graded predictions of $s_{\phi}$ for these smaller constituent steps: $s_{\text{target}}\leftarrow\text{stopgrad}(s_{t}+s_{t+d})/2$, where $s_{t}$ and $s_{t+d}$ are velocities for the small steps and $d$ here represents $\Delta t_{\text{small}}$. This bootstrapping mechanism enables the policy to learn how to make larger, yet accurate, discrete jumps. The parameter $(d,t)$ in Algorithm 3 (line 2) allows sampling various small step sizes $d$ and interpolation times $t$ for this consistency training.
Once trained, this shortcut policy $s_{\phi}$ can generate action sequences with significantly reduced computational cost using a sampling procedure like Algorithm 4 (where $s_{\theta}$ in the algorithm is the trained $s_{\phi}$). The number of steps $M$ in Algorithm 4 can now be set to a small integer (e.g., 1 to 5), drastically reducing inference latency while aiming to maintain high generation quality due to the specialized training of $s_{\phi}$.
1:  Input: Dataset $D$, model $s_{\theta}$, batch size $B$, parameter $p$, market observations dataset $O$
2:  repeat
3:     Sample batch $\{x_{1}^{i}\}_{i=1}^{B}\sim D$, $\{x_{0}^{i}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$, $\{(d^{i},t^{i})\}_{i=1}^{B}\sim p(d,t)$, $\{O_{t}^{i}\}_{i=1}^{B}\sim O$
4:     for $i=1$ to $\lfloor p\cdot B\rfloor$ do
5:        $x_{t}^{i}\leftarrow(1-t^{i})\cdot x_{0}^{i}+t^{i}\cdot x_{1}^{i}$
6:        $s_{\text{target}}^{i}\leftarrow(x_{1}^{i}-x_{0}^{i})$
7:        $d^{i}\leftarrow 0$
8:     end for
9:     for $i=\lfloor p\cdot B\rfloor+1$ to $B$ do
10:        $x_{t}^{i}\leftarrow(1-t^{i})\cdot x_{0}^{i}+t^{i}\cdot x_{1}^{i}$
11:        $s_{t}^{i}\leftarrow s_{\theta}(x_{t}^{i},t^{i},d^{i}|O_{t}^{i})$Add market observation conditioning}
12:        $x_{t+d}^{i}\leftarrow x_{t}^{i}+d^{i}\cdot s_{t}^{i}$Follow ODE trajectory}
13:        $s_{t+d}^{i}\leftarrow s_{\theta}(x_{t+d}^{i},t^{i}+d^{i},d^{i}|O_{t}^{i})$
14:        $s_{\text{target}}^{i}\leftarrow\text{stopgrad}((s_{t}^{i}+s_{t+d}^{i})/2)$Modified stopgrad position}
15:     end for
16:     Update $\theta$ by minimizing $\frac{1}{B}\sum_{i=1}^{B}\|s_{\theta}(x_{t}^{i},t^{i},d^{i}|O_{t}^{i})-s_{\text{target}}^{i}\|^{2}$Use consistent step size}
17:  until convergence
Algorithm 3 Shortcut Policy Training
1:  Input: Market observation $O_{t}$, trained model $s_{\theta}$, steps $M$
2:  Output: Trading action sequence $a_{M}$
3:  Sample $a_{0}\sim\mathcal{N}(0,I)$
4:  $d\leftarrow 1/M$Keep step size consistent with training}
5:  for $k=0$ to $M-1$ do
6:     $t_{k}\leftarrow k\cdot d$
7:     $a_{k+1}\leftarrow a_{k}+d\cdot s_{\theta}(a_{k},t_{k},d|O_{t})$Use same step size d}
8:  end for
9:  return  $a_{M}$ as the generated trading action sequence
Algorithm 4 Shortcut Policy Inference
3.3 Fine-tune Pretrained Model
In the first component, a flow matching policy learns to imitate the strategies of various experts. So it is a general, pre-trained policy. The second component of our framework aims to fine-tune this pre-trained policy to surpass the performance of experts. To enhance the framework’s real-time adaptability to immediate market dynamics, we apply a linear transformation to the actions proposed by the pre-trained model. The parameters of this transformation are rapidly calibrated using the most current market information. This process is more computationally lightweight and efficient compared to recalibrating the traditional HFT models. The output of this calibrated linear transformation is then combined with the initial action from the pre-trained model to yield the final, refined trading decision.
Specifically, this fine-tuning uses a linear transformation of the form $(a\cdot\text{output}+b$) to the action sequence output by the pretrained model. Specifically, if the pretrained model outputs a sequence of planned actions $({a_{t+1},\dots,a_{t+T_{\text{pred}}}})$ over the prediction horizon, where each action $(a_{i})$ is a vector representing the bid and ask offsets for time $(t+i)$ (e.g., $(a_{i}=[\delta_{\text{bid},i},\delta_{\text{ask},i}])$), the fine-tuned action sequence $({a^{\prime}{t+1},\dots,a^{\prime}{t+T_{\text{pred}}}})$ is computed as $a^{\prime}i=a\cdot a_{i}+b)$ for each step $(i)$ from 1 to $(T{\text{pred}})$. Here, $(a)$ is a scalar scaling factor applied element-wise to each vector $(a_{i})$, and $(b)$ is an additive vector offset applied to each vector $(a_{i})$. Both the scalar $(a)$ and the vector $(b)$ (with dimensions matching the action vector, e.g., 2 for bid/ask offsets) are calibrated using a validation set. This adjusted action sequence $(a^{\prime})$ is then used by the policy (typically by executing the first action $(a^{\prime}_{t+1})$) in the environment.
4 Generation of Imitation Learning Data
4.1 Market Simulation and Expert Models
FlowHFT is an imitation learning framework. Therefore, it is crucial to prepare high-quality learning material. In this section, we firstly detail how to generate a set of various market scenarios. Secondly, for each individual scenario, we evaluate a pool of candidate experts—including established traditional algorithms and reinforcement learning agents—and designate the one exhibiting the best performance as the ’expert’ for that context. We use the actions generated by the expert’s strategy to create state-action pairs for training FlowHFT.
We model the mid-price $S_{t}$ as a jump-diffusion process merton1976option , capturing both continuous price movements and sudden jumps. The stochastic differential equation (SDE) governing the mid-price is given by: $dS_{t}=S_{t^{-}}\left(\mu dt+\sigma dB_{H}(t)\right)+S_{t^{-}}(e^{J}-1)dN_{t}$ , where $\mu$ represents the drift coefficient, signifying the expected return of the asset. The asset’s volatility is denoted by $\sigma$, which quantifies the fluctuations in its price. The term $dB_{H}(t)$ corresponds to the increment of a fractional brownian motion process, capturing the continuous component of price movement. Furthermore, $J$ denotes the size of the jump, following a normal distribution $N(\mu_{J},\sigma_{J}^{2})$, which specifies the mean and variance of the jump. Lastly, $dN_{t}$ is a Poisson process with intensity $\lambda_{J}$, modeling the occurrence of jumps, thereby reflecting sudden shifts in the asset’s price.
Next we discuss simulation of order arrival. Order arrivals are modeled using mutually exciting Hawkes processes hawkes1971point ; bacry2015hawkes , effectively capturing the self-exciting and the cross-exciting nature of financial markets. ’self-exciting’ effects indicate that prior buy (or sell) orders increase subsequent arrivals of the same order type, respectively. ’Cross-exciting’ effects describe how buy orders influence sell order arrivals and vice-versa.
The intensity of buy and sell orders, $\lambda_{b}(t)$ and $\lambda_{a}(t)$ respectively, are defined as: $\lambda_{a}(t)=\mu_{a}+\sum_{t_{i}\in\mathcal{N}_{a}}\alpha_{aa}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{b}}\alpha_{ab}e^{-\beta(t-t_{j})}$ and $\lambda_{b}(t)=\mu_{b}+\sum_{t_{i}\in\mathcal{N}_{b}}\alpha_{bb}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{a}}\alpha_{ba}e^{-\beta(t-t_{j})}$ where, $\mu_{b}$ and $\mu_{a}$ are the baseline intensities for buy and sell orders, respectively. The parameters $\alpha_{bb}$ and $\alpha_{aa}$ describe the self-exciting effects of buy and buy, and sell and sell orders, respectively, while $\alpha_{ba}$ and $\alpha_{ab}$ represent the cross-exciting effects between buy and sell orders. The parameter $\beta$ denotes the decay rate, which influences how quickly the effects of past events diminish over time. Finally, $N_{b}$ and $N_{a}$ are the sets containing the timestamps of past buy and sell order arrivals, which are used to model the timing of these events.
To generate diverse expert demonstrations for FlowHFT, we utilized a set of expert candidates, including: (1) the Avellaneda-Stoikov (AS) model avellaneda2008high , a foundational stochastic control framework for optimal quoting that considers inventory risk and market volatility; (2) the Guéant-Lehalle-Fernandez-Tapia (GLFT) model gueant2013dealing , which extends such optimal control approaches, often incorporating aspects such as order flow dynamics; (3) a modified GLFT model incorporating a price drift component, designed to adapt strategies in markets exhibiting directional trends; and (4)a model-free Reinforcement Learning (RL) agent trained with Proximal Policy Optimization (PPO) schulman2017proximal , which learns a policy through direct interaction or simulation to maximize a specified reward function. The detailed mathematical formulations, assumptions, and parameterization specifics for each of these expert models are provided in Appendix [Your Appendix Letter/Number].
4.2 Generation of Market Scenarios
We create market scenarios with varying levels of liquidity—categorized as high, medium, and low. Furthermore, to test the model performance under stress, we introduce scenarios featuring sudden market changes and periods of significantly increased volatility, allowing us to examine how effectively each model responds to dynamic and challenging environments. We simulate different market environments by adjusting the following Simulation parameters:
  1. 1.
    Hurst Component Levels: The Hurst component is determined by the Hurst exponent H, which quantifies the long-term memory and self-similar characteristics of the mid-price time series. A large $H>0.5$ indicates trend-following behavior, a small $H<0.5$ indicates mean-reversion, and $H=0.5$ signifies Markov property of memorylessness.
  2. 2.
    Volatility Levels: Controlled by the volatility parameter $\sigma$ in the mid-price dynamics. High volatility corresponds to larger $\sigma$, inducing more significant price fluctuations.
  3. 3.
    Liquidity Levels: Adjusted through the baseline intensities $\mu_{b}$, $\mu_{a}$, and the excitation parameters in the Hawkes processes. High liquidity corresponds to high arrival rates and frequent orders, while low liquidity corresponds to low arrival rates and infrequent orders.
  4. 4.
    Sudden Market Changes: Introduced by incorporating jump events in the mid-price dynamics with higher intensity $\lambda_{J}$ and larger jump sizes $J$.
In the model, the time step sizes are denoted by $dt$ and are within the range 0.01, 0.02. The drifts, represented by $\mu$, vary among 0.01, 0.05, 0.2. The volatilities, $\sigma$, are considered at values 0.1, 0.2, 0.4. The Hurst Exponent, $H$, is evaluated at 0.3, 0.5, 0.7, indicating the level of long-term memory of the time series. The jump intensities, $\lambda_{j}$, are set at 0.05, 0.1, 0.2. Baseline intensities for buying and selling, $\mu_{a}$ and $\mu_{b}$, are equal and chosen from 10, 20, 40. The strategy employed is the "Dominant Stochastic Strategy," which guides the decision-making process within this stochastic framework. Combining these parameters yields $2\times 3\times 3\times 3\times 3\times 3\times 1=486$ parameter combinations.
4.3 Generation of State-Action Pairs
We generate state-action pairs for each market generated. For each parameter combination, we simulate $100$ episodes. Each episode consists of $1/dt$ time steps (e.g., $100$ steps if $dt=0.01$), each timestamp yields a state-action pair from the expert strategy, here we collect the corresponding state-action pairs from the dominant stochastic strategy among our candidates(AS, GLFT, GLFT-Drift, PPO). In total, we have gathered 1.62 million state-action pairs. These generated state-action pairs serve as the training dataset for the diffusion policy, providing a broad coverage of market conditions and ensuring robust policy learning. This dataset serves as the training data for our diffusion policy framework.
5 Experiments
To evaluate FlowHFT’s adaptability and performance, our experiments were conducted under out-of-sample market conditions. This was achieved by systematically controlling market parameters in the test environments to ensure they differed from those encountered during the training phase. The research questions are: (1) Can FlowHFT effectively generalize strategies learned from expert demonstrations to novel market conditions? (2) Does the integrated fine-tuning mechanism demonstrably improve the performance of actions initially proposed by the pre-trained model? (3) Can the FlowHFT framework achieve improved profitability in these challenging out-of-sample situations?
5.1 Test Market Environment Setting
We set the underlying price process’s memory characteristics by setting the Hurst exponent (H) to values of 0.2, 0.5, and 0.8. A Hurst exponent of H=0.2 generates strongly mean-reverting market conditions, where price changes tend to rapidly reverse direction. Setting H=0.5 simulates a random walk process typical of efficient markets with no memory effect, serving as a crucial baseline. On the other hand, H=0.8 creates strongly persistent or trending markets, where price movements are more likely to continue in their current direction. Second, we introduce variations in the overall market directionality by specifying the drift parameter ($\mu$) as either 0, representing a market with no inherent long-term upward or downward trend, or 0.2, introducing a positive drift to simulate a generally bullish market sentiment. These parameter choices result in 2×3=6 distinct core market scenarios based on trend and memory properties. Within each of these six scenarios, we further differentiate the market microstructure by creating four specific situations based on combinations of market volatility (Vol) and order arrival rate (AR): High Vol/High AR (HH), High Vol/Low AR (HL), Low Vol/High AR (LH), and Low Vol/Low AR (LL). This subdivision is designed to capture a range of typical market states: ’HH’ represents active, potentially news-driven volatile markets; ’HL’ mimics risky, potentially gappy conditions with low liquidity; ’LH’ reflects stable, liquid markets ideal for high-frequency trading; and ’LL’ simulates quiet, perhaps stagnant market periods. This comprehensive matrix of 6×4=24 scenarios provides a rigorous testbed for evaluating the adaptability and performance of HFT strategies across diverse and representative market conditions.
To implement the fine-tuning component, we determine the necessary adjustment parameter using a specific portion of the test market data. Each test period is divided into a validation set and a subsequent trading phase. The validation set period is utilized to calibrate the fine-tuning parameter. The calculated parameter is then used during the trading phase to adjust or modulate the actions proposed by the original pre-trained flow matching model.
5.2 Evaluation Metrics
  • Profit and Loss (PnL): Measures the total profits generated by the trading strategy.
  • Sharpe Ratio (SR): Assesses the risk-adjusted return.
  • Maximum Drawdown (MDD): Evaluates the risk by measuring the largest peak-to-trough decline in PnL.
5.3 Results Analysis
Looking at the results presented from Table 1 to Table 6, a consistent pattern emerges across the various simulated scenarios. Firstly, the GLFT model typically yields superior results compared to the AS model. Critically, our pre-trained flow matching policy model demonstrates performance levels comparable to the GLFT instructor. This finding is significant as it indicates that the pre-trained model successfully learned and internalized effective strategies from its instructors, even within these challenging stochastic environments.
Examining the performance results corresponding to the fine-tuned Flow Matching policy, a clear and significant trend emerges. This adapted model consistently and significantly outperforms both the traditional baseline models (AS and GLFT) and the initial pre-trained model across the tested environments. This finding is paramount, as it validates the second core objective of our framework: we not only demonstrate the ability to effectively learn from expert instructors but also showcase the capacity to surpass their performance. This success is attributed to the fine-tuning process, which empowers the model to "sense" the prevailing market conditions via the validation set calibration and subsequently adjust its behavior, specializing its strategy for the specific market environment encountered.
Table 1: Parameters for Market Conditions: Hurst Exponent $H=0.5$, Drift Rate $\mu=0$, volatility ($\sigma$: {0.02, 0.25}) and arrival rate ($\mu$: {25, 50}). The linear finetuning (ax+b) are the following: HH: action+[0.15,0.24], HL: action+[0.1,0.2], LH: action+[-0.1,0.1], LL: action+[-0.3,0.3],, these parameters can be chosen from an interval, just pick a point as representation The evaluation metrics are tested out of sample, with each methods are tested 1 million trails to obtain significance and convergence.
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
White Noise 1.99 0.06 28.49 0.99 0.04 19.24 2.10 0.31 2.71 1.08 0.22 1.87
AS 24.22 0.09 241.65 13.54 0.09 125.78 25.20 1.05 7.66 13.67 0.72 6.61
GLFT 25.10 0.37 60.57 13.56 0.24 52.55 25.87 1.17 6.95 13.91 0.78 6.14
GLFT-drift 25.10 0.37 60.57 13.56 0.24 52.55 25.87 1.17 6.95 13.91 0.78 6.14
RL-PPO 14.76 0.10 133.61 9.29 0.08 103.85 26.74 0.81 10.13 19.80 0.46 14.56
Pretrained Flow Policy 21.99 0.34 56.94 13.28 0.23 53.43 22.29 1.5 4.07 12.10 0.91 4.2
Finetuned Flow Policy 25.82 0.38 58.38 14.43 0.26 49.64 26.79 1.62 3.62 22.31 0.96 5.17
Table 2: Performance Comparison Under Different Conditions (Hurst Exponent $H=0.5$, Drift Rate $\mu=0.02$).Same measures as the above table. Flow Policy select GLFT-drift strategy as the expert to learn from. The linear finetuning (ax+b) for each action is HH: action+[0.1,0.25], HL: action+[0.1,0.2], LH: action+[-0.3,0.3], LL: action+[-0.4,0.7], these parameters can be chosen from multiple choices, just pick one as representation
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
WhiteNoise 1.99 0.06 28.90 0.99 0.04 19.51 2.10 0.28 2.87 1.09 0.19 1.98
AS 24.16 0.08 245.49 13.51 0.09 127.70 25.18 0.92 8.02 13.61 0.63 6.98
GLFT 25.11 0.36 61.39 13.55 0.23 53.26 25.21 0.99 7.45 13.62 0.66 6.59
GLFT-drift 25.87 0.37 61.26 14.15 0.24 53.21 40.82 1.10 7.69 26.44 0.87 6.97
RL-PPO 30.74 0.16 186.1 20.47 0.14 138.88 24.79 0.72 10.58 32.49 0.65 15.28
Pretrained Flow Policy 22.72 0.35 57.09 13.28 0.23 53.43 21.90 1.64 3.97 11.96 0.98 4.05
Finetuned Flow Policy 27.45 0.39 61.31 15.15 0.26 50.95 42.81 1.17 8.44 27.12 0.93 7.42
Tables 4-7 expand the volatility and arrival rate options to investigate a broader spectrum of unseen scenarios. (Volatility, $\sigma$: {0.05, 0.2}) and arrival rates, $\mu$: {30, 60})
Table 3: Performance Comparison Under Different Conditions (Hurst Exponent $H=0.2$, Drift Rate $\mu=0.0$). Same measures as the above table. Diffusion Policy select GLFT-drift strategy as the expert to learn from. The linear finetuning (ax+b) for each action is HH: action+[0.02,0.2], HL: action+[0.05,0.15], LH: action+[-0.15,0.15], LL: action+[-0.1,0.1], these parameters can be chosen from multiple choices, just pick one as representation
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
WhiteNoise 2.41 0.07 56.73 1.27 0.05 39.31 2.49 0.24 14.23 1.31 0.17 9.94
AS 28.82 0.14 346.8 16.23 0.13 216.69 29.28 0.78 48.86 16.16 0.54 40.22
GLFT 29.16 0.36 134.33 15.78 0.23 115.22 29.29 0.92 41.34 16.1 0.61 35.2
GLFT-drift 29.16 0.36 134.33 15.78 0.23 115.22 29.29 0.92 41.34 16.1 0.61 35.2
Pretrained Flow Policy 26.61 0.35 119.73 14.76 0.22 112.78 26.56 1.16 26.76 14.61 0.71 26.58
Finetune Flow Policy 29.72 0.36 136.52 15.80 0.23 114.62 31.43 1.06 36.25 16.82 0.75 29.33
Table 4: Performance Comparison Under Different Conditions (Hurst Exponent $H=0.2$, Drift Rate $\mu=0.02$).Same measures as the above table. Diffusion Policy select GLFT-drift strategy as the expert to learn from. The linear finetuning (ax+b) for each action is HH: action+[-0.1,0.1], HL: action+[-0.1,0.1], LH: action+[-0.2,0.2], LL: action+[-0.3,0.3], these parameters can be chosen from multiple choices, just pick one as representation
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
WhiteNoise 2.41 0.07 57.51 1.28 0.05 39.86 2.49 0.22 14.47 1.32 0.16 10.11
AS 28.78 0.14 352.12 16.22 0.13 219.9 29.26 0.73 49.74 16.15 0.5 40.93
GLFT 29.14 0.35 136.07 15.77 0.23 116.71 29.27 0.86 42.06 16.1 0.57 35.81
GLFT-drift 29.96 0.36 136.4 16.53 0.24 117.15 35.23 0.96 44.55 21.2 0.69 38.11
Pretrained Flow Policy 26.67 0.35 121.2 14.9 0.21 114.29 26.5 1.11 27.2 14.52 0.67 27.05
Finetuned Flow Policy 31.3 0.36 140.71 18.41 0.24 127.55 38.6 1.14 42.58 28.07 0.75 49.98
Table 5: Performance Comparison Under Different Conditions (Hurst Exponent $H=0.8$, Drift Rate $\mu=0.0$).Same measures as the above table. Diffusion Policy select GLFT-drift strategy as the expert to learn from. The linear finetuning (ax+b) for each action is HH: action+[-0.02, 0.1], HL: action+[0.1, 0.1], LH: action+[-0.1, 0.1], LL: action+[-0.1, 0.1], these parameters can be chosen from multiple choices, just pick one as representation
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
WhiteNoise 2.48 0.10 11.72 1.31 0.07 8.13 2.51 0.28 3.27 1.32 0.20 2.32
AS 27.89 0.17 71.05 15.88 0.16 44.13 29.17 0.91 8.29 16.10 0.63 7.55
GLFT 29.00 0.57 21.05 15.77 0.36 19.88 29.21 1.10 6.63 16.07 0.73 6.38
GLFT-drift 29.00 0.57 21.05 15.77 0.36 19.88 29.21 1.10 6.63 16.07 0.73 6.38
Pretrained Flow Policy 26.33 0.67 17.37 14.53 0.36 19.26 26.61 1.66 3.80 14.45 0.93 4.57
Finetuned Flow Policy 30.31 0.67 17.77 17.61 0.39 19.66 29.59 1.50 3.42 16.65 0.96 4.22
Table 6: Performance Comparison Under Different Conditions (Hurst Exponent $H=0.8$, Drift Rate $\mu=0.02$).Same measures as the above table. Diffusion Policy select GLFT-drift strategy as the expert to learn from. The linear finetuning (ax+b) for each action is HH: action+[-0.1, 0.1], HL: action+[-0.1, 0.1], LH: action+[-0.1, 0.1], LL: action+[-0.1, 0.1], these parameters can be chosen from multiple choices, just pick one as representation
High Volatility & High Demand High Volatility & Low Demand Low Volatility & High Demand Low Volatility & Low Demand
PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$ PnL $\uparrow$ SR $\uparrow$ MDD $\downarrow$
WhiteNoise 2.48 0.10 11.92 1.32 0.07 8.27 2.51 0.26 3.42 1.33 0.18 2.42
AS 27.83 0.17 72.41 15.86 0.15 44.95 29.15 0.83 8.70 16.09 0.58 7.92
GLFT 28.98 0.56 21.46 15.76 0.35 20.23 29.19 1.00 6.95 16.07 0.67 6.68
GLFT-drift 30.05 0.57 21.08 16.74 0.37 19.94 35.22 1.11 5.71 21.24 0.80 5.60
Pretrained Flow Policy 26.64 0.66 17.57 14.52 0.34 19.47 26.74 1.61 3.79 14.50 0.87 4.68
Finetuned Flow Policy 33.17 0.62 19.64 19.20 0.38 20.61 32.57 1.61 3.01 19.02 1.01 3.80
6 Conclusion
In conclusion, we propose a novel imitation learning framrwork FlowHFT, the first application of flow matching policy in a financial stochastic control task. We have confirmed its capability to learn effective trading strategies in stochastic environments. A key strength of FlowHFT is its ability to integrate knowledge from expert strategies across diverse market scenarios into a single model, enabling prompt adaptation to prevailing conditions. Furthermore, the incorporated grid search mechanism successfully refines actions, especially in situations where those experts were suboptimal. FlowHFT thus represents a promising advancement in developing adaptive and high-performing HFT strategies.
Appendix A Expert Market-Making Models
A.1 Avellaneda-Stoikov (AS) Model
The Avellaneda-Stoikov (AS) model [2] optimizes bid and ask spreads by maximizing the expected utility of terminal wealth. Taking the coefficients of drift, jump intensity, self-excitation and mutual-excitation to be zero in above settings , The value function $u(s,x,q,t)$ is governed by the Hamilton-Jacobi-Bellman (HJB) equation (details in Appendix):
$\frac{\partial u}{\partial t}+\frac{1}{2}\sigma^{2}\frac{\partial^{2}u}{\partial s^{2}}$ $+\max_{\delta_{b}}\lambda_{b}\left(\delta_{b}\right)\left[u\left(s,x-s+\delta_{b},q+1,t\right)-u(s,x,q,t)\right]$
$+\max_{\delta_{a}}\lambda_{a}\left(\delta_{a}\right)\left[u\left(s,x+s+\delta_{a},q-1,t\right)-u(s,x,q,t)\right]=0,$ (5)
with terminal condition $u(s,x,q,T)=-e^{-\gamma\left(x+qS_{T}\right)}$.
The optimal bid and ask spreads are:
$\delta_{\text{bid}}^{\text{AS}}$ $=\frac{\gamma\sigma^{2}\tau}{2}+\frac{1}{\gamma}\ln\left(1+\frac{\gamma}{k}\right)-q\gamma\sigma^{2}\tau,$ (6)
$\delta_{\text{ask}}^{\text{AS}}$ $=\frac{\gamma\sigma^{2}\tau}{2}+\frac{1}{\gamma}\ln\left(1+\frac{\gamma}{k}\right)+q\gamma\sigma^{2}\tau,$ (7)
In the model, $\gamma$ is defined as the risk aversion parameter, which quantifies the trader’s preference for avoiding risk relative to seeking returns. The variable $\sigma$ represents the asset’s volatility, indicating the degree of price fluctuation over time. The term $\tau=T-t$ denotes the remaining time horizon, calculating the time left until the end of the trading period. The parameter $k$ measures the sensitivity of order arrivals to price spreads, demonstrating how spread changes influence trading activity. Lastly, $q$ is the current inventory level, reflecting the amount of the asset held by the trader.
A.2 Guéant-Lehalle-Fernandez-Tapia (GLFT) Model
The Guéant-Lehalle-Fernandez-Tapia (GLFT) model [7] offers an approximate solution for optimal spreads by incorporating additional terms into the HJB equation. The assumptions are keeping the same as AS model expect. There is no limitation to the terminal time T. The optimal spreads are:
$c_{1}$ $=\frac{1}{\gamma}\ln\left(1+\frac{\gamma}{k}\right),$ (8)
$c_{2}$ $=\sqrt{\frac{\gamma}{2Ak}\left(1+\frac{\gamma}{k}\right)^{\left(\frac{k}{\gamma}+1\right)}},$ (9)
$\delta_{\text{bid}}^{\text{GLFT}}$ $=c_{1}+\frac{\sigma c_{2}}{2}+\sigma c_{2}q,$ (10)
$\delta_{\text{ask}}^{\text{GLFT}}$ $=c_{1}+\frac{\sigma c_{2}}{2}-\sigma c_{2}q,$ (11)
In the model, $A$ denotes the base arrival rate of orders. The coefficients $c_{1}$ and $c_{2}$ are used to adjust the bid and ask spreads based on factors such as risk aversion, volatility, and time horizon. Lastly, $q$ represents the current inventory level, indicating the amount of the asset that is held.
A.3 Modified Guéant-Lehalle-Fernandez-Tapia (GLFT) Model with Drift
Keeping everything unchanged except for the drift part, the new modified GLFT solution which hasn’t been shown in the original paper[7].
$|q|=Q_{\max}$
$dS_{t}=\mu dt+\sigma dW_{t}$
$(\delta^{b}_{\infty})^{*}=\frac{1}{\gamma}\ln(1+\frac{\gamma}{k})+\left[-\frac{\mu}{\gamma\sigma^{2}}+\frac{2q+1}{2}\right]\sqrt{\frac{\sigma^{2}}{2kA}(1+\frac{\gamma}{k})^{1+\frac{k}{\gamma}}}$
$(\delta^{a}_{\infty})^{*}=\frac{1}{\gamma}\ln(1+\frac{\gamma}{k})+\left[\frac{\mu}{\gamma\sigma^{2}}-\frac{2q-1}{2}\right]\sqrt{\frac{\sigma^{2}}{2kA}(1+\frac{\gamma}{k})^{1+\frac{k}{\gamma}}}$
A.4 Proximal Policy Optimization (PPO) for High-Frequency Market Making
To apply Proximal Policy Optimization (PPO) in a high-frequency trading (HFT) environment, we formulate the trading process as a Markov Decision Process (MDP). The input to the policy model consists of high-dimensional market state features, such as limit order book snapshots, recent trade volumes, price trends, and microstructure indicators sampled at millisecond intervals. The action space can include discrete actions like placing a buy/sell/cancel order at a specific price level. The reward function is designed to capture profit-and-loss (PnL) adjusted for market impact and inventory risk, including components like realized spread, execution cost, and inventory penalties. The objective is to train a policy that maximizes expected cumulative rewards.
In the context of market making, the RL framework can be formalized as a Markov Decision Process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where, $S$ represents the state space, and $A$ denotes the action space. The state transition probabilities are symbolized by $P$, which dictates the likelihood of transitioning from one state to another given a particular action. $R$ is the reward function, which assigns a reward based on the state and action taken. Finally, $\gamma\in[0,1]$ is the discount factor, used to determine the present value of future rewards, allowing for the evaluation of long-term versus immediate benefits.
For the space state, at each discrete time step $t$, the state $s_{t}$ encapsulates essential information required for decision-making:
$s_{t}=(t,x_{t},q_{t},S_{t},B_{t-1}-A_{t-1}),$
In the model, $t$ is the timestamp indicating the specific point in time under consideration. The variable $x_{t}$ denotes the current cash position, while $q_{t}$ represents the current inventory, which refers to the number of shares held. The term $S_{t}$ is the current underlying stock price, providing a snapshot of the stock’s market value at time $t$. $A_{t-1}-B_{t-1}$ is the difference between the ask and bid quotes from the previous time step, reflecting the spread.
For the action space, the agent’s action $a_{t}$ consists of setting new ask and bid quotes relative to the mid-price $M_{t}$. Specifically:
$a_{t}=(\delta^{\text{ask}}_{t},\delta^{\text{bid}}_{t}),$
In the model, $\delta_{\text{ask}t}$ represents the offset of the ask quote from the mid-price, which sets the ask price to $M_{t}+\delta{\text{ask}t}$. Conversely, $\delta{\text{bid}t}$ is the offset of the bid quote from the mid-price, setting the bid price to $M_{t}-\delta{\text{bid}_{t}}$.
For the reward function, it is designed to balance profitability from the bid-ask spread against the risk associated with inventory holding. The formulation is:
$r_{t}=\text{PnL}_{t}-\lambda q_{t}^{2},$
$PnL_{t}$ represents the profit and loss at time $t$, derived from executed trades. The parameter $\lambda$ is a risk aversion parameter that penalizes large inventory positions.
After defining the MDP, policy gradient methods is used to optimize the policy $\pi_{\theta}(a|s)$ parameterized by $\theta$ to maximize the expected cumulative reward:
$J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$
where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T})$ denotes a trajectory sampled from the policy. The output of the policy network is a parameterized probability distribution over actions, from which the actual trading action is sampled. The derivation related to policy gradient is in Appendix.